15 students at UDM are selected randomly. 5 are Nursing majors, 5 are BioChem majors, and 5 are Economics majors. If two are selected, what is the probability, without replacement, that the first would be a Nursing major and the second would be BioChem?
Answer #1

Here total number of students are 15

Now we are suppose to select two students

First one is nurse and for that probability is \frac{5}{15}=0.3333

Now second one is BioChem for that probability is \frac{5}{14} =0.3571

So required probability is 0.3333*0.3571 = 0.1190